### v5

# DeepVideo-R1: Video Reinforcement Fine-Tuning via Difficulty-aware Regressive GRPO

Abstract
Recent works have demonstrated the effectiveness of reinforcement learning (RL)-based post-training for enhancing the reasoning capabilities of large language models (LLMs). In particular, Group Relative Policy Optimization (GRPO) has shown impressive success using a PPO-style reinforcement learning algorithm with group-normalized rewards. However, the effectiveness of GRPO in Video Large Language Models (VideoLLMs) remains underexplored. In this paper, we explore GRPO and identify two issues that hinder effective learning: (1) reliance on safeguards, and (2) vanishing advantage. To mitigate these challenges, we propose DeepVideo-R1, a video large language model trained with Reg-GRPO (Regressive GRPO) and difficulty-aware data augmentation. Reg-GRPO reformulates the GRPO loss function as a regression task that directly predicts the advantage in GRPO, eliminating the need for safeguards such as clipping and min operations. This directly aligns the model with the advantages, providing guidance to prefer better outputs. The difficulty-aware data augmentation strategy augments input prompts/videos to target solvable difficulty levels, enabling diverse reward signals. Our experimental results show that our approach significantly improves video reasoning performance across multiple benchmarks.
1 Introduction
Large Language Models (LLMs) [radford2021learning, touvron2023llama, achiam2023gpt, team2023gemini] have demonstrated remarkable abilities in understanding, reasoning, and generating text across diverse domains. Their success stems from next-token prediction over vast corpora, which enables the emergence of complex reasoning patterns and world knowledge. Building on this progress, recent research has extended LLMs into the video domain, giving rise to Video Large Language Models (VideoLLMs) [wang2024qwen2, wang2024internvideo2, bai2025qwen2, zhu2025internvl3, zhang2025videollama]. These models aim to unify video understanding and language generation, enabling capabilities such as temporal event reasoning, video question answering, and video-to-text summarization.
Despite their rapid evolution, current VideoLLMs still struggle with complex reasoning tasks, which require temporal, spatial, and semantic understanding over video sequences. Since standard supervised fine-tuning fits instruction data rather than reasoning processes, it is limited in improving reasoning capabilities. To address this, reinforcement learning (RL)-based post-training [zhang2024direct, ahn2025isr] has emerged as a compelling paradigm. RL provides a mechanism to optimize models beyond likelihood objectives, aligning them with reward signals that encode human preference or task-specific success. Recently, Group Relative Policy Optimization (GRPO) [shao2024deepseekmath, guo2025deepseek] has shown promise by using group-based advantages and relative preference signals to enhance reasoning capabilities.
While GRPO has achieved strong results in text-based tasks, its application to VideoLLMs remains underexplored. In this work, we investigate the application of GRPO to VideoLLMs and identify two key limitations that hinder effective training: (1) reliance on stabilizers such as minimum and clipping operations, which often suppress gradients and impede convergence, and (2) the vanishing advantage problem, where extremely easy or difficult samples yield zero advantages, thereby removing the training signal.
To overcome these limitations, we propose DeepVideo-R1, a video large language model trained with two key innovations: Regressive GRPO (Reg-GRPO) and difficulty-aware data augmentation. Reg-GRPO reformulates the GRPO objective as a regression problem that directly predicts group-based advantage values. This simple yet effective reformulation enables direct alignment between model outputs and the advantage values, eliminating the need for stabilizers while ensuring stable convergence. We also introduce a difficulty-aware augmentation that dynamically adjusts the difficulty of video-text inputs. For easy samples, we perturb the video content to inject uncertainty; for hard samples, we provide auxiliary reasoning cues. This strategy diversifies the reward landscape, mitigating the vanishing advantage problem and promoting balanced learning across difficulty levels.
Figure 1: DeepVideo-R1 significantly improves the reasoning capabilities of VideoLLMs. Our VideoLLM, DeepVideo-R1, is trained to explicitly predict the advantage $\hat{A}^{(i)}$ through Regressive GRPO loss. Notably, model training becomes significantly effective and achieves a 10.1 performance improvement compared to GRPO.
Our experimental results demonstrate the effectiveness of DeepVideo-R1 on multiple challenging video reasoning benchmarks such as SEED-Bench-R1, LongVideoBench, and NExTGQA, achieving superior performance over recent video LLMs such as Qwen2.5-VL [bai2025qwen2] (Figure 1). Notably, our model achieves consistent improvements on both in-distribution and out-of-distribution tasks, indicating robust generalization capabilities. These results underscore the benefits of combining a regression-based RL objective with data augmentation for training large-scale multimodal reasoning models.
Our main contributions are listed as:
  • We introduce Reg-GRPO, a novel optimization scheme that casts GRPO as a regression task over group-based advantage values, eliminating heuristic stabilizers such as clipping and min operations, and mitigating the vanishing gradient issue.
  • We develop a difficulty-aware augmentation framework that modulates video-text inputs with adaptive difficulty scaling, video cue injection, and noise perturbation to generate richer and more effective reward signals.
  • We propose DeepVideo-R1, a video large language model trained with two key innovations: Regressive GRPO (Reg-GRPO) and difficulty-aware data augmentation. Our experimental results demonstrate that DeepVideo-R1 significantly improves the reasoning capabilities of VideoLLMs on complex video reasoning tasks.
2 Related Work
Video Large Language Models (VideoLLMs).
Large Language Models (LLMs) [brown2020language, yang2024qwen2, grattafiori2024llama] have exhibited strong generalization and reasoning capabilities across a wide range of domains, including knowledge-intensive tasks [jin2025search, luo2025graph, park2023graph, park2024generative], mathematical reasoning [shao2024deepseekmath, luo2023wizardmath], and scientific domains [zheng2025large, bai2025intern, park2024llamo]. Building on these capabilities, Video Large Language Models (VideoLLMs) have extended LLM reasoning to dynamic video domains, achieving notable performance across tasks [wu2024longvideobench, fu2024video, li2024mvbench, shen2024longvu], such as video question answering [bai2025qwen2, zhu2025internvl3, zhang2025videollama, li2023videochat, ko2023large, li2024llava] and video captioning [xu2023mplug, cheng2024videollama, zhang2024video]. Despite their impressive performance, VideoLLMs remain limited on long video inputs and fine-grained video understanding tasks that require detailed spatiotemporal reasoning [gao2017tall, krishna2017dense, xiao2024can]. Most existing methods primarily emphasize video perception or short-context understanding, often relying on static supervised fine-tuning objectives that fail to capture reasoning dynamics. To address these challenges, we leverage a reinforcement learning-based fine-tuning approach to improve the reasoning and generalization capabilities of VideoLLMs.
RL-based fine-tuning.
Multiple works [shao2024deepseekmath, guo2025deepseek, jaech2024openai, team2025kimi, lee2025vidchain, lee2025captioning] have significantly improved the reasoning capabilities of LLMs through reinforcement learning (RL), including DPO [rafailov2023direct] and RLHF [ouyang2022training]. Recently, variants of RL-based fine-tuning [zhu2024self, gao2024rebel] have explored direct reward-regression losses derived from RL objectives. A key development in this direction is Group Relative Policy Optimization (GRPO), an RL algorithm proposed in [shao2024deepseekmath] that computes group-wise normalized rewards to stabilize training and improve efficiency. Motivated by GRPO, several approaches have demonstrated substantial improvements in the reasoning abilities of MLLMs across image [liu2025visual, yang2025r1, zhan2025vision, shen2025vlm, huang2025vision, huang2025boosting, liu2025noisyrollout, wang2025vl] and video tasks [feng2025video, li2025videochat, chen2025exploring, wang2025timezero, wu2025st, tan2025reason]. While existing approaches [li2025videochat, wang2025timezero, chen2025exploring] have primarily focused on defining appropriate reward functions tailored to each visual task, some concurrent works [huang2025boosting, liu2025noisyrollout] instead study practical issues that arise during GRPO training, aiming to further enhance model reasoning. In this work, we propose a learning algorithm that directly regresses advantages instead of increasing the likelihood of high-advantage responses. Additionally, we design difficulty-aware data augmentation to provide diverse and dense learning signals.
3 Methods
In this section, we present a video large language model named DeepVideo-R1, which is trained with Regressive GRPO (Reg-GRPO) and difficulty-aware data augmentation for effective video context reasoning. We first introduce post-training methods for VideoLLMs, such as proximal policy optimization and group-relative policy optimization (GRPO), and discuss the limitations of GRPO: reliance on heuristic safeguards and vanishing advantage. Then, we propose Reg-GRPO, which improves the RL-based GRPO by transforming the objective into a simpler yet more effective regression loss, eliminating heuristic safeguards such as the min and clipping operations. Finally, we present a novel difficulty-aware data augmentation, which alleviates the vanishing advantage problem by modulating the difficulty of samples.
3.1 RL-based Fine-Tuning
Proximal Policy Optimization (PPO) [wang2020truly]
is one of the most widely used actor-critic RL algorithms for fine-tuning (video) large language models. For example, RLHF [huang2024n+] applies the PPO algorithm. Given the input sample $\boldsymbol{x}$, PPO optimizes the model $\pi_{\theta}$ with the following objective:
$\begin{split}&\mathcal{L}_{\text{PPO}}\left(\theta\right)=-\mathbb{E}_{\boldsymbol{x},\;\boldsymbol{y}\sim\pi_{\theta_{\text{old}}}\left(\cdot|\boldsymbol{x}\right)}\\ &\frac{1}{\left\lvert\boldsymbol{y}\right\rvert}\sum_{t=1}^{\left\lvert\boldsymbol{y}\right\rvert}\min\left[\frac{\pi_{\theta}\left(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t}\right)}{\pi_{\theta_{\text{old}}}\left(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t}\right)}{A}_{t},\text{clip}\left(\frac{\pi_{\theta}\left(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t}\right)}{\pi_{\theta_{\text{old}}}\left(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t}\right)},1-\epsilon,1+\epsilon\right)A_{t}\right],\end{split}$ (1)
where $\boldsymbol{y}$ is sampled from the policy model $\pi_{\theta}$, $\epsilon$ is a hyperparameter, $\pi_{\theta_{\text{old}}}$ is the old model and the advantage $A_{t}$ is calculated with generalized advantage estimation (GAE) [schulman2015high] using rewards and a trained value function $V_{\psi}$. Although the PPO algorithm aligns human preferences with the policy model outputs effectively, it requires substantial memory and computational resources since the value function is typically another model comparable in size to the policy model.
Group Relative Policy Optimization (GRPO) [shao2024deepseekmath]
addresses the problem of PPO [wang2020truly] by approximating the learnable value function with the average reward of multiple sampled outputs. Concretely, given an input sample $\boldsymbol{x}$, the model samples multiple output sequences $\left\{\boldsymbol{y}^{(i)}\right\}_{i=1}^{G}$ from the old policy model $\pi_{\theta_{\text{old}}}$ and then trains the policy model $\pi_{\theta}$ with the following objective:
$\begin{split}&\mathcal{L}_{\text{GRPO}}\left(\theta\right)=\mathbb{E}_{\boldsymbol{x},\;\left\{\boldsymbol{y}^{(i)}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}\left(\cdot|\boldsymbol{x}\right)}}\\ &\frac{1}{\left\lvert\boldsymbol{y}^{(i)}\right\rvert}\sum_{t=1}^{\left\lvert\boldsymbol{y}^{(i)}\right\rvert}\Biggl\{\min\left[\frac{\pi_{\theta}\left(y_{t}^{(i)}|\boldsymbol{x},\boldsymbol{y}_{<t}^{(i)}\right)}{\pi_{\theta_{\text{old}}}\left(y_{t}^{(i)}|\boldsymbol{x},\boldsymbol{y}_{<t}^{(i)}\right)}\hat{A}^{(i)}_{t},\text{clip}\left(\frac{\pi_{\theta}\left(y_{t}^{(i)}|\boldsymbol{x},\boldsymbol{y}_{<t}^{(i)}\right)}{\pi_{\theta_{\text{old}}}\left(y_{t}^{(i)}|\boldsymbol{x},\boldsymbol{y}_{<t}^{(i)}\right)},1-\epsilon,1+\epsilon\right)\hat{A}^{(i)}_{t}\right]\\ &\quad\quad\quad\quad\quad-\beta\mathcal{D}_{\text{KL}}\left[\pi_{\theta}||\pi_{\text{ref}}\right]\Biggl\},\end{split}$ (2)
where $\beta$ corresponds to a hyperparameter and $\mathcal{D}_{\text{KL}}$ is the KL-divergence. Here, $\hat{A}^{(i)}$ is advantage calculated based on the relative reward within the group, which is formulated as $\hat{A}^{(i)}=\frac{\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)-\mu_{r}}{\sigma_{r}}$ where $\mu_{r},\sigma_{r}$ denotes the average and standard deviation values of a set of rewards in the group, respectively. Although GRPO has shown its success, GRPO has two limitations that hinder the effective model optimization: reliance on heuristic constraints and vanishing advantage problems.
Reliance on safeguards.
GRPO optimizes the model with safeguards implemented using the min and clipping functions to avoid extreme changes in the model. However, the PPO-style clipping function induces zero gradient for samples where the value of $\pi_{\theta}\left(\boldsymbol{y}|\boldsymbol{x}\right)$} is too different from the value of $\pi_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)$}. It cannot guarantee that the model $\pi_{\theta}\left(\boldsymbol{y}|\boldsymbol{x}\right)$ stays close to $\pi_{\theta_{\text{ref}}}$ if it is already far from $\pi_{\theta_{\text{ref}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)$ [hsu2020revisiting]. Similarly, GRPO also suffers from this phenomenon due to the PPO-style hard constraints, and it deteriorates the effective model training. The analysis in [yu2025dapo] also shows that an upper clipping threshold restricts the probability increase of the “exploration” token. This indicates that the safeguards in GRPO negatively influence model optimization.
Vanishing advantage problem.
The vanishing advantage problem [wang2025vl] indicates that the advantage of each sample within the group becomes zero, when the rewards of outputs in the group are equal. It is problematic since the model cannot receive any signals from the training sample where the advantage is zero for every response. In particular, we observe that this issue often arises when training samples are either too easy or too difficult for the current model. Training samples with extreme difficulty levels show worse performance than those with moderate difficulty levels.
3.2 Regressive GRPO
Here, we present a Reg-GRPO (Regressive Group Relative Policy Optimization), which reformulates GRPO into the regression task, removing safeguards such as the min and clipping functions. This reformulation enables the model to directly predict the advantages, resulting in improved alignment of the model with the preference. Following existing RL-based works [peters2007reinforcement, peters2010relative, rafailov2023direct, gao2024rebel], the Reg-GRPO loss function is derived from the RL objective that maximizes the expected reward with the KL constraints between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$.
RL objective.
The objective of our reinforcement learning algorithm for each iteration is to maximize rewards while preventing $\pi_{\theta}$ from making excessive changes relative to $\pi_{\theta_{\text{old}}}$:
$\pi_{\theta}^{*}=\underset{\pi_{\theta}}{\mbox{arg}\max}\ \mathbb{E}_{\boldsymbol{x},\boldsymbol{y}\sim\pi_{\theta}\left(\cdot|\boldsymbol{x}\right)}\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)-\lambda\;\mathbb{E}_{\boldsymbol{x}}\left[\mathbb{D}_{\text{KL}}\left(\pi_{\theta}\left(\cdot|\boldsymbol{x}\right)||\pi_{{\theta_{\text{old}}}}\left(\cdot|\boldsymbol{x}\right)\right)\right],$ (3)
where $\lambda$ is the hyperparameter that adjusts the strength of the KL-divergence. Following prior works [rafailov2023direct, gao2024rebel], the closed-form solution to the above equation (Eq. (3)) can be obtained by minimum relative entropy problem:
${\pi}_{\theta}^{*}\left(\boldsymbol{y}|\boldsymbol{x}\right)=\frac{1}{Z\left(\boldsymbol{x}\right)}\pi_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)\exp\left(\frac{1}{\lambda}\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)\right),\forall\boldsymbol{x},\boldsymbol{y}$ (4)
where $Z\left(\boldsymbol{x}\right)=\sum_{\boldsymbol{y}}{\pi}_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)\exp\left(\frac{1}{\lambda}\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)\right)$ is a partition function. However, since calculating the partition function $Z\left(\boldsymbol{x}\right)$ is expensive, it is hard to obtain $\pi_{\theta}^{*}$ exactly.
Reg-GRPO Loss.
To address these issues, we propose Reg-GRPO (Regressive GRPO) loss, which learns the policy model to regress the advantage $\hat{A}^{(i)}$ using the reparameterization, removing the normalization term $Z\left(\boldsymbol{x}\right)$. Specifically, the advantage for the $i$-th sample is defined as
$\hat{A}^{(i)}=\frac{\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)-\mu_{r}}{\sigma_{r}},$ (5)
where $\mu_{r},\sigma_{r}$ denote the average and standard deviation values of a set of rewards in the group, respectively. We can also rewrite Eq. (4) to express the reward $\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)$ in terms of the optimal model $\pi_{\theta}^{*}$, which can be formulated as
$\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)=\lambda\cdot\left(\log Z\left(\boldsymbol{x}\right)+\log\left(\frac{{\pi}_{\theta}^{*}\left(\boldsymbol{y}|\boldsymbol{x}\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)}\right)\right)\quad\forall\boldsymbol{x},\boldsymbol{y}.$ (6)
Since the reward can be expressed through the optimal policy $\pi_{\theta}^{*}$ (Eq. (6)), the advantage can be equivalently written as $\hat{A}^{(i)}=\frac{\rho^{*}\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)-\mu_{\rho^{*}}}{\sigma_{\rho^{*}}}$, where $\rho^{*}\left(\boldsymbol{x},\boldsymbol{y}\right)$ is defined as $\rho^{*}\left(\boldsymbol{x},\boldsymbol{y}\right)=\log\frac{\pi_{\theta}^{*}\left(\boldsymbol{y}|\boldsymbol{x}\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)}$ and $\mu_{\rho^{*}},\sigma_{\rho^{*}}$ denote mean and standard deviation of $\left\{\rho^{*}\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)\right\}_{i=1}^{G}$, respectively. Interestingly, we can see that $Z\left(\boldsymbol{x}\right)$ is naturally removed during the reformulation.
Building on this insight, we define the predictive advantage, which estimates the advantage calculated by normalizing rewards within a group of samples, using the current policy $\pi_{\theta}$ as
$\hat{A}^{(i)}_{\theta}=\frac{\rho\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)-\mu_{\rho}}{\sigma_{\rho}},\quad\rho\left(\boldsymbol{x},\boldsymbol{y}\right)=\log\frac{\pi_{\theta}\left(\boldsymbol{y}|\boldsymbol{x}\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{y}|\boldsymbol{x}\right)},$ (7)
where $\mu_{\rho},\sigma_{\rho}$ are mean and standard deviation of $\left\{\rho\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)\right\}_{i=1}^{G}$, respectively. Then, we optimize the policy by minimizing the gap between the target advantage $\hat{A}$ and its predicted counterpart $\hat{A}_{\theta}$ using Reg-GRPO (Regressive GRPO), which is defined as:
$\mathcal{L}_{\text{Reg-GRPO}}\left(\theta\right)=\mathbb{E}_{\boldsymbol{x},\left\{\boldsymbol{y}^{(i)}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}\left(\cdot|\boldsymbol{x}\right)}\left\{\left(\hat{A}^{(i)}-\hat{A}_{\theta}^{(i)}\right)^{2}-\beta\;\mathbb{D}_{\text{KL}}\left[\pi_{\theta}||\pi_{\text{ref}}\right]\right\},$ (8)
Similar to GRPO, we regularize the update with the KL divergence. The proposed Reg-GRPO loss serves as an effective alternative for optimizing group-level objectives, showing better performance than GRPO. In our experiments, we demonstrate that this formulation leads to faster convergence and improved policy quality. The detailed derivation procedure of Reg-GRPO is in Appendix LABEL:app_sec:derivation.
Figure 2: Overview of the difficulty-aware data augmentation. First, we assess the difficulty of responses given the input video and question using Eq. (9). For hard samples, it augments the input prompts with the reasoning cues extracted from successful reasoning paths (Difficulty-decreasing augmentation), while the easy samples are perturbed with the noise (Difficulty-increasing augmentation). The scale of the guidance level or noise level is adaptively determined based on the difficulty of the current sample.
3.3 Difficulty-aware data augmentation
In this section, we present a difficulty-aware data augmentation framework, which addresses the vanishing advantage problem in GRPO. This issue arises when training samples are either too easy or too difficult, leading to uniform rewards across multiple responses. As a result, the advantage values become zero, erasing the learning signal. Our augmentation strategy mitigates this issue by modulating the difficulty of inputs to increase variance in predicted rewards, thereby preserving informative gradients for effective model optimization.
Specifically, given an input sample $\boldsymbol{x}=\left(\boldsymbol{v},\boldsymbol{q}\right)$, where $\boldsymbol{v}$ and $\boldsymbol{q}$ denote a video and a question, respectively, we first generate multiple responses and then compute the average reward $\frac{1}{G}\sum_{i=1}^{G}\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}^{(i)}\right)$ for the sample. We then measure the difficulty of $\boldsymbol{x}$ by comparing its average reward with the average reward of samples in a replay buffer $\mathbb{B}^{W}$, which consists of samples $\boldsymbol{x}_{\text{rep}}$ and their corresponding outputs $\left\{\boldsymbol{y}^{(i)}_{\text{rep}}\right\}_{i=1}^{G}$ from the most recent $W$ steps. Formally, the difficulty $\Delta_{\mathcal{R}}\left(\boldsymbol{x}\right)$ of the input sample $\boldsymbol{x}$ is calculated as:
$\Delta_{\mathcal{R}}\left(\boldsymbol{x}\right)=\mathbb{E}_{\left(\boldsymbol{x}_{\text{rep}},\left\{\boldsymbol{y}_{\text{rep}}^{(i)}\right\}_{i=1}^{G}\right)\in\mathbb{B}^{W}}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{R}\left(\boldsymbol{x}_{\text{rep}},\boldsymbol{y}_{\text{rep}}^{(i)}\right)\right]-\frac{1}{G}\sum_{j=1}^{G}\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}^{(j)}\right).$ (9)
Instead of using only the sample reward $\mathcal{R}\left(\boldsymbol{x},\boldsymbol{y}\right)$, we use the average reward of samples in the replay buffer as a reference value to account for model evolution.
Building on the difficulty metric in Eq. (9), we adaptively adjust each training sample to balance the learning signal. For samples identified as too easy (i.e., high reward and low difficulty), we increase task complexity by perturbing the input—for example, by adding Gaussian noise to video inputs or masking video frames—to encourage the model to attend to more informative cues. Conversely, for too difficult samples (i.e., low reward and high difficulty), we inject auxiliary reasoning hints or visual cues that simplify the temporal context, helping the model focus on core reasoning paths instead of failing due to overwhelming difficulty. This dynamic modulation leads to an appropriate difficulty distribution across training, preventing the collapse of advantage values while ensuring that each update provides a meaningful gradient signal.
Difficulty-decreasing augmentation.
For difficult samples ($\Delta_{\mathcal{R}}\left(\boldsymbol{x}\right)>0$), we ease the difficulty of the sample by providing auxiliary reasoning cues that guide the model toward generating correct reasoning. Concretely, we first augment the input prompt $\boldsymbol{q}$ with the ground-truth answer and generate multiple reasoning trajectories using the VideoLLM. Among these reasoning trajectories, we select the response $\boldsymbol{y}_{\text{gt}}$ with the highest reward and extract a partial reasoning trace of it. This trace is then incorporated into the original prompt to form a modified prompt $\tilde{\boldsymbol{q}}$ containing structured hints that guide the model’s reasoning toward the correct solution. To maintain adaptive control, the guidance level is scaled according to the sample’s difficulty magnitude. Harder samples receive stronger reasoning cues, while moderately difficult samples are given lighter guidance. By adaptively providing guidance for challenging inputs, this augmentation mitigates vanishing gradients in difficult cases and facilitates more stable convergence through progressively refined reasoning.
Difficulty-increasing augmentation.
Conversely, for easy samples ($\Delta_{\mathcal{R}}\left(\boldsymbol{x}\right)<0$), we employ a difficulty-increasing augmentation to enhance task complexity and encourage the model to explore more diverse reasoning trajectories. We perturb the visual input $\boldsymbol{v}$ to create a harder input $\tilde{\boldsymbol{v}}$ by introducing frame-level Gaussian noise, thereby slightly degrading perceptual fidelity while preserving overall semantic structure. The intensity of the noise is proportionally scaled by the difficulty magnitude, ensuring that easier samples receive stronger perturbations and moderately easy samples remain stable. This adaptive corruption expands the diversity of generated reasoning trajectories. By inducing distributed rewards, the augmentation ensures that even trivial samples provide informative gradients, avoiding zero learning signals throughout the optimization process.
Table 1: Performance on SEED-Bench-R1 validation split and LongVideoBench. In-Dist. means in-distribution dataset.